# Towards Open-World Gesture Recognition

### v1

@@ 4 Large-Scale Data Collection
    Procedure: Participants wore a watch-like device with an IMU sensor on their non-dominant hand to collect data. During each session, participants were prompted to perform the target gesture through a chime or vibration, while an experimenter recorded the start and stop times of each gesture. To ensure an accurate representation of real user behavior, randomization was implemented throughout the study. A total of 100 sessions were conducted, with each session including all participants performing 50 instances of a single gesture recurrence for each of the four gestures. These gestures were paired with with 25 contexts, and each of the contexts mimics a real-life context. Examples of the 25 contexts are: standing/walking with hand up to the chest level/ with hand hanging, riding in a car a passenger, pushing stroller/cart with the other hand, holding a cup or something else in the other hand, sitting at the desk with elbow on the desk and hand in the air/ with arm laying on the desk/ with the elbow on the arm rest and hand in the air/ with the arm on an arm rest, laying down on back/side, lounging on the sofa (horizontal/slouched posture) with hand laying on the sofa or laps, cuddling/arm around someone else, stationary biking, walking up and down stairs, jogging, leaning over, picking something up. The contexts are also differentiated by if the participant is looking at the device or not. We also collected negative (i.e., non-gesture) examples under various activities such as jogging, running, biking, clapping, driving, waving, cleaning, etc. These examples are similar to the negative examples from [63]. This methodology aimed to capture the natural variability in users’ motions without introducing any repetition bias.
@@ 5 Design Engineering Approach
Conventional machine learning approaches optimize the recognition model to best fit the data once. In an OWGR process illustrated in Figure 1, we additionally must optimize the policies by which the recognition model continually updates itself. In the open-world gesture recognition process, five stages are outlined: device deployment, data logging, data curation, retraining trigger, and updating policy. The final two stages highlight the two crucial design considerations integral to an OWGR process that allows offline analysis:
@@
    When to update the recognition model? How do we define the start and stop of a batch, and how frequently should we perform model updates?
@@
    How to update the recognition model? Which continual learning method provides the ideal tradeoff between plasticity and stability while minimizing compute and memory requirements?
Therefore, we propose a design engineering approach to perform a systematic evaluation of an OWGR process without long-term deployment (including user studies over multiple sessions, model deployment and run-time optimization, continuous data logging and collection etc.). Furthermore, evaluation of an OWGR process involves system-level parameters beyond the recognition model hyper-parameters.
@@ 5.1 Functional Design and Function Carriers of OWGR
In this paper, we focus on the conceptual design of the OWGR system, following the methodology in [25]. Initially, we introduce OWGR as a functional design, identifying its main function (Open-World Gesture Recognition) and key sub-functions (Data Logging, Data Curation, Retraining Trigger, Updating Policy, Device Deployment), along with their interrelationships, shown in Figure 1.
Later, we establish solution principles, specifically for the Updating Policy function. We evaluate potential function carriers, considering controllable and uncontrollable parameters, and perform envelope analyses. This helps us understand the system requirements for these carriers to effectively fulfill their main function.
Our study mainly examines two functions, excluding Data Logging, Data Curation and Device Deployment, which require dedicated online analysis studies. This approach aids in quantitative parameter investigation, enhancing our understanding of continual learning methods for open-world gesture recognition. A comprehensive functional design is beyond this paper’s scope.
In this paper, we study six function carriers. This function carrier determines which continual learning method to use to update the gesture recognition model. By varying this function carrier, we can explore the design question: how to update the recognition model. We evaluated the following five different continual learning methods, and a baseline, across the aforementioned families in Section 2:
@@ 5.2 Surrogate Task Model
In envelope analysis, a surrogate model is a simplified variant of a complex, computationally heavy model. It involves identifying and varying controllable and uncontrollable system parameters for system evaluation. Controllable parameters can be fine-tuned for optimization, while uncontrollable ones help predict potential performance variations (sensitivity analysis) [25]. Envelope analysis is performed by altering one parameter at a time, with other parameters set to default.
We propose a surrogate task model in which we can vary different task settings. The task setting in each case determines the specific definition of each task which is represented by the task descriptor $t_{i}$ in a triplet $(x_{i},y_{i},t_{i})$. By varying this setting, we can answer the other design question: when to update the recognition model, because the task setting defines the start and stop of a batch, the size of a batch, and the frequency of model updates, etc. We have the following parameters in the surrogate task model that determines the task setting:
@@
    granularity of tasks: The tasks can be either fine-grained or coarse. For example, in the case of new context, coarse task granularity is standing and fine task granularity is standing with hand up to the chest level. On the other hand, in the case of new gesture, a coarse task granularity contains multiple gestures for the model to learn at one time and fine task granularity contains fewer gestures. This parameter is controllable.
@@
  3. 3.
    new gesture: We include the granularity of task, which is represented by number of gestures in one task. The more gestures a model needs to learn per task, the coarser the granularity of a task is. In addition to one non-gesture (null) class, we vary the number of gesture classes between one and three. We also include order of tasks. The default settings are three target gestures and a random ordering of tasks.
@@ 5.3 Evaluation Metrics
We use accuracy and forgetting as the evaluation measures for the OWGR model, with definitions adopted from [7] as follows:
@@
We report on accuracy and forgetting of the final model, obtained by evaluating each task after learning the entire task sequence.
Figure 4: QuartzNet architecture for our gesture recognition model.
Figure 5: Effect of Task Setting on Performance for New Context Each sub-figure represents one task setting. Each box plot shows mean, median, and quartile data. a) Order of Contexts: Change of context order does not significantly affect average accuracy and forgetting measure for most methods. The X-axis represents a specific order of the tasks. E-H is easy-to-hard, H-E is hard-to-easy. b) Granularity of Contexts: Some methods perform particularly better or worse in fine-grained context than coarse context. The X-axis represents that if the context is coarse or fine. c) Number of Contexts: Different methods exhibits various performances when number of contexts (tasks) increases. The X-axis represents the total number of contexts (tasks).
Figure 6: Effect of Task Order on Performance for New Gesture. The order of tasks does not have a significant effect on the accuracy and forgetting metrics for most methods except finetuning. Each box plot shows mean, median, and quartile data. The X-axis represents the order of the tasks in New Gesture case, H-E is hard-to-easy, E-H is easy-to-hard.
Figure 7: Effect of Number of Gestures Per Task on Performance for New Gesture. A larger number of gestures per task leads to a significant decrease in accuracy for all methods. Each box plot shows mean, median, and quartile data. The X-axis represents the number of gestures in one group (task) in new gesture case. In each task, there is a non-gesture representing background activities with no gesture performed. This non-gesture is not counted in the number of gestures.
@@ 5.4 Implementation Details
We implement the continual learning algorithms based on an open-source generalizing continual learning framework [10] in PyTorch. For data processing, we use a sliding window approach to segment the data for both training data and testing data preparation, we use a window size of 120 and a window step size of 60. We adopt the same parameters for debouncing thresholds in [49].
For the hyperparameter search of each continual learning method, we conduct an approach that first tries to decay each hyperparameter separately, and decays all if none of these individual decays to achieve accuracy within the finetuning margin. This process is then repeated until the stability decay criterion is met.
For the specific setup of the training framework, we define maximal plasticity search with a coarse learning rate grid $\left\{1e^{-2},5e^{-3},1e^{-3},5e^{-4},1e^{-4}\right\}$. We set the finetuning accuracy drop margin to 0.2 and the decaying factor to 0.5. We use a Stochastic Gradient Descent with a momentum of 0.9 and batch size of 128. We use a max of 100 training epochs with early stopping and annealing of the learning rate. That is the learning rate decays with factor 10 after 10 unimproved iterations of the validation accuracy and the training process should terminate after 15 unimproved iterations of training.
We use a simplified version of QuartzNet architecture [24], a state-of-the-art convolutional architecture for speech recognition, for our gesture recognition model described by $f$ in Equation 1. It is used as the classifier that takes IMU signals as input and predicts among the gesture classes. The selection of the QuartzNet model is based on preliminary experiments. This model achieves state-of-the-art performance in real-time gesture recognition, striking a balance between accuracy and latency. This is attributed to the model’s quantization, which allows it to be compact enough to operate in real-time on smaller devices. To make the model deployable on more memory-constrained devices, we have removed the residual connections, grouped point-wise convolutions, and channel shuffle without scarifying the accuracy performance. The overall model architecture is shown in Figure 4.
@@ 6 Results
This section describes the results from the envelope analysis by varying function carriers’ parameters in each case: new context, new user, and new gesture.
We describe the results of varying the surrogate task model for each of the three use cases:
6.1 New Context
Figure 5 consists of three sub-figures which describe the effect of different task settings separately for the new context case. Prior work suggests that an easy-to-hard task ordering might achieve better performance than a hard-to-easy ordering [58]. However, Figure 5 a) shows that the impact of the task order on the average accuracy and forgetting measure is insignificant. We do observe that the variance of the performance scores is considerably smaller for easy-to-hard ordering.
In Figure 5 b), when setting the granularity of contexts to fine-grained as the task-setting, we see that finetuning shows a lower accuracy and a higher forgetting measure, whereas methods such as LWF, PackNet, and replay reach a higher accuracy while maintaining a low forgetting measure. A coarse context setting has more training data, as a coarse context consists of multiple fine-grained contexts. This suggests that these continual learning methods can successfully optimize plasticity (transfer knowledge of previous tasks to new tasks) to achieve data-efficiency.
Figure 5 c) shows that a lower number of total tasks can stabilize the average performance of the continual learning methods. A larger number of total tasks leads to an increased forgetting measure as the model’s capacity for learning new tasks is limited. However, PackNet and LwF can successfully re-use the previous knowledge to produce an increased accuracy on new tasks. On the other hand, finetuning introduces a significant forgetting measure accompanied by a low accuracy score when the number of tasks is large.
6.2 New Gesture
In Figure 6, we see that in the new gesture case, the ordering of tasks does not significantly impact the performance. This is similar to the new context case. Figure 7 shows that more gestures in the task leads to decreasing accuracy, while the forgetting measure is maintained to a low level except the baseline finetuning (which also exhibits increasing forgetting measure).
6.3 New User
In Figure 8, we observe that a larger number of users increases both accuracy and forgetting metrics with finetuning and replay methods. In contrast, most continual learning methods such as PackNet, MAS and LWF achieve high accuracy while maintaining low forgetting measure with larger number of total users. This suggests that, in the new user case, PackNet, MAS, and LwF can better transfer knowledge of old users to new users.
@@ 7 Discussion
7.1 Revealing Design Guidelines Through Envelope Analysis
We revisit the two general design questions proposed in Section 5 and provide insights from our results.
7.1.1 When to update the recognition model?
@@
This question asks the best task setting for each case, and we use the envelope analysis from the surrogate task model to answer this question. Most of the parameters in the surrogate task model are controllable by the designers. We advise that a coarse context usually gives better results for the new context case. A coarse context is also more practical, as it simplifies the activity recognition needed to produce a task descriptor. For the specific scenario in new gesture case where a user wants to author a new gesture for the same command, our advice is that it is better to use a single gesture classifier for each command as the fewer gestures a model needs to classify, the better. Moreover, our results show that the order of the applications (tasks) does not affect performance. For the new user case, we see the best results when the number of different users is large with LwF and PackNet, as we observe that a larger number of users increases the overall accuracy for these two methods. In practice, a model needs to be pre-trained and it is encouraged to pre-train the model on a large-scale dataset if available.
7.1.2 How to update the recognition model?

### v2

@@ 2.3 Large-Scale Data Collection
    Procedure: Participants wore a watch-like device with an IMU sensor on their non-dominant hand to collect data. During each session, participants were prompted to perform the target gesture through a chime or vibration, while an experimenter recorded the start and stop times of each gesture. To ensure an accurate representation of real user behavior, randomization was implemented throughout the study. A total of 100 sessions were conducted, with each session including all participants performing 50 instances of a single gesture recurrence for each of the four gestures. These gestures were paired with with 25 contexts, and each of the contexts mimics a real-life context. Examples of the 25 contexts are: standing/walking with hand up to the chest level/ with hand hanging, riding in a car a passenger, pushing stroller/cart with the other hand, holding a cup or something else in the other hand, sitting at the desk with elbow on the desk and hand in the air/ with arm laying on the desk/ with the elbow on the arm rest and hand in the air/ with the arm on an arm rest, laying down on back/side, lounging on the sofa (horizontal/slouched posture) with hand laying on the sofa or laps, cuddling/arm around someone else, stationary biking, walking up and down stairs, jogging, leaning over, picking something up. [^2] Table 1 in the Appendix shows the full list of contexts. The contexts are also differentiated by if the participant is looking at the device or not. We also collected negative (i.e., non-gesture) examples under various activities such as jogging, running, biking, clapping, driving, waving, cleaning, etc. These examples are similar to the negative examples from [66]. This methodology aimed to capture the natural variability in users’ motions without introducing any repetition bias.
@@ 3 Design Engineering Approach
Conventional machine learning approaches optimize the recognition model to best fit the data once. In an OWGR process, as illustrated in Figure 1, we additionally must optimize the policies by which the recognition model continually updates itself. In the open-world gesture recognition process, we outline five stages: (1) device deployment; (2) data logging; (3) data curation; (4) retraining trigger; and (5) updating policy. The final two stages highlight the two crucial design considerations that are integral and critical to an OWGR process that allows offline analysis:
@@
    When do we update the recognition model? How do we define the start and stop of a batch, and how frequently should we perform model updates?
@@
    How do we update the recognition model? Which continual learning method provides the ideal trade-off between plasticity and stability, while minimizing processing and memory requirements?
We tackle these two considerations by adapting a design engineering approach [25, 26] to perform a systematic evaluation of an OWGR process without long-term deployment, including user studies over multiple sessions, model deployment and run-time optimization, continuous data logging and collection, and so on. Further, the evaluation of an OWGR process involves system-level parameters beyond the recognition model hyper-parameters.
(a) New Context: Each new task requires one user to perform the same set of gestures under new context, such as standing, walking, riding in a car, laying down, etc. In this context, we have 10 tasks in total, due to size restrictions in presentation, we only show results up to Task 6.
(b) New User: Each new task inquires 6 new users to perform the same set of gestures in the same context. This is akin to a model learning to generalize across different users’ behavior patterns.
Figure 3: We examine the accuracy of a gesture recognition model, in an open-world gesture recognition problem, tested on Task $n$, while the model is trained incrementally from Task $n$ to Task $N$. $N$ is the total number of tasks, and the model is trained by five continual learning methods and one baseline method: Finetuning. For instance, in a subplot with title Task 2, we test the model, progressively trained from Task 2 through Task 10, exclusively on Task 2 and report 9 data points. It lacks the data points from the model trained on Task 1 and tested on Task 2, because the tasks were introduced sequentially; a model trained on Task 1 has not been exposed to Task 2. Consequently, the subplot for Task 2 omits these results. This pattern continues with Task 3 and subsequent tasks, with a subplot of Task $n$ only reports $N-n+1$ data points (per method). The observed deep decline in accuracy from Finetuning in each subplot can be attributed to the model’s tendency to forget earlier tasks as it learns new ones, which is referred to catastrophic forgetting. We proposed five continual learning methods to address this catastrophic forgetting problem in open-world gesture recognition, namely SI, replay, LWF, PackNet and MAS, which are introduced in detail in Section 3.1. In this Figure, we illustrate the results from performing preliminary experiments on two use cases of open-world gesture recognition. We report average accuracy (forgetting) in the legend. Forgetting is the measure of the decrease in accuracy. We discuss these two measures in detail in Section 3.3.
@@ 3.1 Functional Design and Function Carriers of OWGR
Following prior work [25, 26], we begin by identifying the main function (Open-World Gesture Recognition) and its key sub-functions (Data Logging, Data Curation, Retraining Trigger, Updating Policy, Device Deployment), along with their interrelationships.
Having established a function model, we then seek function carriers, or solution principles, and specifically for the Updating Policy function. We evaluate potential function carriers by considering their controllable and uncontrollable parameters through envelope analyses that simulate emergent outcomes as a function of the controllable and uncontrollable parameters in the system. This helps us understand the system requirements for these function carriers to effectively fulfill their functions.
Our study mainly examines two functions, Retraining Trigger and Updating Policy, which both require dedicated online analysis studies. This approach aids in quantitative parameter investigation, thereby enhancing our understanding of continual learning methods for open-world gesture recognition. A comprehensive functional design of the entire system is beyond this paper’s scope.
We here study six function carriers. The study of these function carriers will determine which continual learning method to use to update the gesture recognition model. By varying the function carrier, we can explore the design question: how do we update the recognition model?. We evaluate the following five different continual learning methods, and a baseline, across the aforementioned families in Section 1.4:
@@
Additionally, we carried out a preliminary offline experiment with internal participants who were asked to perform gestures under two distinct scenarios. In this experiment, we tested the above six different continual learning methods, aiming to adapt the gesture recognition model. The outcomes of this study are presented in Figure 3. This preliminary analysis primarily highlights three key findings:
  • Naive fine-tuning with new task data leads to significant forgetting of previously learned tasks, underscoring the necessity for continual learning methods to address this issue. This emphasizes the critical nature of the open-world gesture recognition paradigm.
  • Continual learning methods demonstrate varied effectiveness across different scenarios; a method that shows superior performance in one case may not be the best choice for another. This underlines the importance of conducting comprehensive quantitative analyses on large-scale open-world gesture recognition datasets to provide design guidance.
  • The execution of this preliminary study required substantial resources and time, as participants needed to perform gestures in a variety of contexts and combine different gestures. This underscores the critical need for online experiments to supplement or replace such extensive offline analyses.
@@ 3.2 Surrogate Task Model
In envelope analysis, a surrogate model is a simplified variant of a complex, computationally involved model. It involves identifying and varying controllable and uncontrollable system parameters to evaluate a system with respect to its emerging outcomes. Controllable parameters can be fine-tuned for optimization, while uncontrollable ones help predict potential performance variations (sensitivity analysis) [25]. Envelope analysis is performed by altering one parameter at a time, with the other parameters held constant.
We here propose a surrogate task model in which we can vary different task settings. The task setting in each case determines the specific definition of each task which is represented by the task descriptor $t_{i}$ in a triplet $(x_{i},y_{i},t_{i})$. By varying this setting, we can answer the other design question: when do we update the recognition model?, because the task setting defines the start and stop of a batch, the size of a batch, and the frequency of model updates, and so on. We have the following parameters in the surrogate task model that determines the task setting:
@@
    Granularity of tasks: The tasks can be either fine-grained or coarse. For example, in the case of new context, coarse task granularity is standing and fine task granularity is standing with hand up to the chest level. This parameter is controllable.
@@ 3.3 Evaluation Metrics
We use accuracy and forgetting as the evaluation measures for the OWGR model, with definitions adopted from prior work [7] as follows:
@@
In this paper, we do not report accuracy using different metrics, such as F1 scores, because the contribution is not to propose state-of-the-art gesture recognition models. Instead, the focus is on comparing the relative accuracy and the forgetting measures between different continual learning methods to provide design guidance for developing open-world gesture recognition methods. Furthermore, given the extensive experiments conducted under various settings in this paper, it is impractical to include and report on different metrics. Therefore, we report on the final accuracy and forgetting, obtained by evaluating each task after learning the entire task sequence.
Figure 4: Effect of Task Setting on Performance for New Context Each sub-figure represents one task setting. Each box plot shows mean, median, and quartile data. a) Order of Contexts: Change of context order does not significantly affect average accuracy and forgetting measure for most methods. The X-axis represents a specific order of the tasks. E-H is easy-to-hard, H-E is hard-to-easy. b) Granularity of Contexts: Some methods perform particularly better or worse in fine-grained context than coarse context. The X-axis represents that if the context is coarse or fine. c) Number of Contexts: Different methods exhibits various performances when number of contexts (tasks) increases. The X-axis represents the total number of contexts (tasks)
@@ 3.4 Implementation Details
We implement the continual learning algorithms based on an open-source generalizing continual learning framework [11] in PyTorch. For data processing, we use a sliding window approach to segment the data for both training data and testing data preparation, we use a window size of 120 and a window step size of 60. We adopt the same parameters for debouncing thresholds in [51]. For the hyperparameter search of each continual learning method, we conduct an approach that first tries to decay each hyperparameter separately, and decays all if none of these individual decays to achieve accuracy within the finetuning margin. This process is then repeated until the stability decay criterion is met For the specific setup of the training framework, we define maximal plasticity search with a coarse learning rate grid $\left\{1e^{-2},5e^{-3},1e^{-3},5e^{-4},1e^{-4}\right\}$. We set the finetuning accuracy drop margin to 0.2 and the decaying factor to 0.5. We use a Stochastic Gradient Descent with a momentum of 0.9 and batch size of 128. We use a max of 100 training epochs with early stopping and annealing of the learning rate. That is the learning rate decays with factor 10 after 10 unimproved iterations of the validation accuracy and the training process should terminate after 15 unimproved iterations of training.
We use a simplified version of QuartzNet architecture [24], a state-of-the-art convolutional architecture for speech recognition, for our gesture recognition model described by $f$ in Equation 1. It is used as the classifier that takes IMU signals as input and predicts among the gesture classes. The selection of the QuartzNet model is based on preliminary experiments. We began with a literature review on state-of-the-art classification models for time series data and tested over 10 models. Our experiments revealed that the QuartzNet model achieves state-of-the-art performance in real-time gesture recognition, striking an optimal balance between accuracy and latency. We further improved the performance by hyperparameter sweeping making it suitable for IMU data classification. This is attributed to the model’s quantization, which allows it to be compact enough to operate in real-time on smaller devices. To make the model deployable on more memory-constrained devices, we have removed the residual connections, grouped point-wise convolutions, and channel shuffle without scarifying the accuracy performance.  [^3] Figure 2 in the Appendix illustrates the details of the model architecture.
@@ 4 Results
This section describes the results of varying the surrogate task model for each of the two use cases.
Figure 4 consists of three sub-figures which describe the effect of different task settings separately for the new context case. Prior work suggests that an easy-to-hard task ordering might achieve better performance than a hard-to-easy ordering [61]. However, Figure 4 a) shows that the impact of the task order on the average accuracy and forgetting measure is insignificant. We do observe that the variance of the performance scores is considerably smaller for easy-to-hard ordering. In Figure 4 b), when setting the granularity of contexts to fine-grained as the task-setting, we see that finetuning shows a lower accuracy and a higher forgetting measure, whereas methods such as LWF, PackNet, and replay reach a higher accuracy while maintaining a low forgetting measure. A coarse context setting has more training data, as a coarse context consists of multiple fine-grained contexts. This suggests that these continual learning methods can successfully optimize plasticity (transfer knowledge of previous tasks to new tasks) to achieve data-efficiency. Figure 4 c) shows that a lower number of total tasks can stabilize the average performance of the continual learning methods. A larger number of total tasks leads to an increased forgetting measure as the model’s capacity for learning new tasks is limited. However, PackNet and LwF can successfully re-use the previous knowledge to produce an increased accuracy on new tasks. On the other hand, finetuning introduces a significant forgetting measure accompanied by a low accuracy score when the number of tasks is large.
For the new user case, we observe from Figure 5 that a larger number of users increases both accuracy and forgetting metrics with finetuning and replay methods. In contrast, most continual learning methods such as PackNet, MAS and LWF achieve high accuracy while maintaining low forgetting measure with larger number of total users. This suggests that, in the new user case, PackNet, MAS, and LwF can better transfer knowledge of old users to new users.
@@ 5 Discussion
5.1 Balance between Stability and Plasticity
Continual learning aims to strike a balance between two key objectives: stability and plasticity. Stability refers to the ability to retain previously learned knowledge over time, avoiding catastrophic forgetting. Plasticity, on the other hand, refers to the ability to acquire new knowledge from new data distributions, avoiding overfitting to old data. Stability is typically measured by evaluating the degree of forgetting or performance degradation on previously learned tasks or data distributions. Plasticity is measured by assessing the model’s accuracy or performance on new, unseen data distributions. When designing OWGR processes, it is crucial to consider this stability-plasticity trade-off. Design choices and techniques should be evaluated based on their ability to achieve an appropriate balance between these two objectives.
5.2 Design Guidelines from Envelope Analysis
5.2.1 When do we update the recognition model?
@@
This question asks for the best task setting for each case. This involves defining the boundaries between tasks and determining when a new task has commenced. Once a new task is detected, a retraining trigger (Figure 2) activates the update of the recognition model using a continual learning method. Thus, we use the results from the envelope analysis from the surrogate task model to answer this question, and we explore which task settings yield the most favorable outcomes. Most of the parameters in the surrogate task model are controllable by the designers. We advise that a coarse context usually gives better results for the new context case. A coarse context is also more practical, as it simplifies the activity recognition needed to produce a task descriptor. The accuracy of a task descriptor determines the granularity of tasks; a less accurate descriptor results in a coarser context, while a more precise descriptor leads to a finer-grained context. We previously discussed that the new context case may require classification of the current activity context, which can be performed by AR/VR HMDs and wearables via activity detection through different sensors [10, 31]. The precision of state-of-the-art activity classifiers, acting as task descriptors, is sufficiently high that it does not adversely affect the primary gesture recognition model. Therefore, these task descriptors are generalizable descriptors and do not require re-training with new real-world data. For the new user case, we see the best results when the number of different users is large with LwF and PackNet, as we observe that a larger number of users increases the overall accuracy for these two methods. In practice, a model needs to be pre-trained and it is encouraged to pre-train the model on a large-scale dataset if available.
5.2.2 How do we update the recognition model?